# Trajectory-Dependent Electronic Energy Losses in Ion Range Simulations

### v1

@@ I Introduction
Ion irradiation of materials is encountered in a wide range of research and industrial applications, including semiconductor doping [1], ion-beam lithography [2], and in the study of radiation-induced damage in nuclear materials [3]. Understanding and predicting the effects of ion irradiation is crucial for optimizing material modification processes and for assessing irradiation-induced material damage [4]. Computational modeling provides a powerful and cost-effective approach for simulating ion-matter interactions and insight into these radiation effects.
@@
The predictive capability of computational models depends on the accurate description of interactions between energetic ions and target materials. These interactions lead to a reduction in the kinetic energy of the ion, quantified by the stopping power: $\frac{dE}{dx}=S_{e}+S_{n}$, where $S_{e}$ is the electronic stopping power (ESP) representing the energy transfer from the energetic ion to target electrons; and $S_{n}$ is the nuclear stopping power (NSP), which arises from interactions between energetic ions and target atomic nuclei [5].
@@
Ion irradiation processes can be modeled using molecular dynamics (MD) simulations or Monte Carlo (MC) methods based on the binary collision approximation (BCA). MD simulations dynamically resolve atomic-scale interactions through Newtonian mechanics, enabling high-accuracy modeling of NSP using interatomic potentials [6, 7]. In contrast, MC-based BCA methods model ion interactions and energy losses statistically, sacrificing atomic-scale accuracy in favor of computational efficiency, allowing for the simulation of many irradiation events. MDRANGE is an MD-based simulation code that was specifically developed to overcome the computational challenge of using MD to predict ion range profiles [8], and offers an efficiency comparable to that of BCA methods with the atomic-scale accuracy of MD.
@@
The ESP is widely modeled as a scalar velocity-dependent retarding force according to Lindhard theory [9], which has been expanded to improve accuracy across different energy regimes. Values for the ESP may be obtained numerically or from computational programs such as PSTAR for protons [10], MSTAR for heavier ions (3Li to 18Ar) [11], or SRIM (The Stopping and Range of Ions in Matter) for all ions. [12] These programs combine theoretical frameworks, scaling laws, and empirical data to provide predictions for the average ESP of ions in materials.
@@
However, projectile ions can travel along trajectories in which they experience lower than average atomic and electronic densities. This effect is known as channeling and occurs when the ions are directed by interatomic potentials along the principal crystal axes (axial) or between planes or rows of atoms (planar) [9, 13]. In these channeling geometries, projectile dynamics are affected less by lattice nuclei, and they lose their kinetic energy mainly through electronic interactions. Ions can also experience varying electronic interactions in these geometries because of the lower electron densities they encounter. These effects lead to a reduction in the total stopping power on average [14] and result in higher ion ranges (penetration depths) in channeling, compared to non-channeling, geometries [13].
@@
Real-time time-dependent density functional theory (rt-TDDFT) has been utilized to simulate the temporal evolution of quantum electronic interactions of energetic ions in materials [15, 16, 17, 18, 6, 19, 20]. While rt-TDDFT can provide an accurate, first-principles description of electronic energy losses, its computational cost is prohibitive for large-scale ion irradiation simulations. In the approach outlined by Correa et al. [21], a trajectory- and energy-dependent scalar ESP can be obtained by analyzing the ion’s rate of energy loss as a function of its displacement in short trajectories and thus utilized in larger-scale simulations.
@@
Recently, Tamm et al. [22] developed a model that uses rt-TDDFT energy loss data to parametrize a coupling function capturing the effects of electronic interactions on a projectile ion, dependent on its local environment, in a two-temperature molecular dynamics (TTMD) framework. This approach enables modeling of a trajectory-dependent ESP with ab initio accuracy in large-scale ion irradiation simulations. The model is implemented for the MD code LAMMPS [23] in the USER-EPH plugin [24].
@@
In this work, we present the implementation of the model developed by Tamm et al. [22], which we hereby refer to as the unified two-temperature model (UTTM), in the MD-based code MDRANGE. This new implementation allows modeling ion ranges over micrometers in depth with ab initio based trajectory-dependent electronic energy losses, with statistics from tens of thousands of ions, which is inaccessible by full MD. Through comparison to other models in suitably chosen geometries, and to experimental measurements where such are available, this method also provides a tool for validating the parametrized coupling functions of the UTTM model. We investigate the performance of the UTTM model implementation in MDRANGE through simulations of ion range profiles in a variety of materials along both channeling and non-channeling trajectories.
@@ II Electronic Stopping Power Models
For any ESP model, the simulated ion range profiles vary along different crystal orientations, due to the different atomic environments encountered [13]. Adjusting the ESP model further modifies these predictions. In this work, we compare trajectory-dependent ion range profile predictions in a variety of materials using different models for the ESP. We categorize two distinct ESP model types: (1) velocity-dependent models for which values can be obtained using SRIM or rt-TDDFT, that apply a scalar retarding force to the projectile ion dependent on its instantaneous velocity, without accounting for the ion’s local environment, and (2) trajectory-dependent models, which explicitly consider the local ion environment in the calculation of the retarding force, i.e. the UTTM model.
II.1 Velocity-Dependent Electronic Stopping
@@
For range simulations along random directions, we employ the velocity-dependent ESP calculated by SRIM [12], which is widely utilized in radiation damage and ion range studies. The ESP is calculated by SRIM semi-empirically, and is based on theoretical models fit to experimental data where it is available. This experimental data is scarce, particularly for heavier ions at low velocities. In these cases, the calculated ESP is extrapolated to the low-velocity regime, introducing uncertainties in modeling the electronic energy losses of slow ions [25].
@@
Moreover, a significant limitation of SRIM is that it does not account for the ordered lattice or electronic structure of materials. Instead, the ESP calculated by SRIM is generally valid for random ion trajectories in target materials. For trajectories where the local electronic and nuclear environments differ significantly from the average random trajectory, such as along channeling directions, electronic energy losses are not accurately modeled [26].
@@
Unlike the semi-empirical approach, rt-TDDFT directly resolves the energy exchanged between projectile ions and the target electrons, which allows for a velocity- and trajectory-dependent description of electronic energy losses. A scalar ESP force on the ion can be calculated for chosen trajectories as outlined by Correa et al. [21]. While this provides higher fidelity in modeling the ESP of ions which travel along channels, it does not account for the local ion environment. Consequently, it does not accurately model electronic energy losses for projectiles which are either not significantly channeled, or have been scattered out of the channel. Nevertheless, the electronic energy losses for strongly channeled ions are captured correctly in range simulations employing a constant rt-TDDFT value for the ESP. Hence, this approach yields accurate predictions of the maximum ion range $R_{max}$ [26], which is typically determined in ion range experiments [27].
II.2 Trajectory-Dependent Electronic Stopping
In order to correctly capture the electronic energy losses for any given ion trajectory, a model that explicitly accounts for the local electronic environment of the ion in the calculation of the ESP is necessary. We implemented the UTTM model from Tamm et al. [22] in MDRANGE to incorporate trajectory-dependent electronic energy losses in ion range simulations.
@@ II.2 Trajectory-Dependent Electronic Stopping
The UTTM model is based on Langevin dynamics, wherein particles in a system experience frictional forces and corresponding random forces via the fluctuation-dissipation theorem. The UTTM model modifies the traditional Langevin dynamics by replacing forces on individual atoms with correlated many-body forces, and a tensorial frictional force describes the ESP acting on the atoms [22]. The force on an atom is defined by:
@@
where $\alpha=\alpha(\rho)$ is the coupling function which is dependent on the electron density, $\rho$, experienced by the ion at any position. The correlated random forces are determined by a set of independent white noise Gaussian variables $\langle\xi_{I}(t)\rangle_{I}$ given by:
@@
where $T_{e}$ is the temperature of the electronic system [22]. The $\alpha(\rho)$ coupling component is determined from ab initio rt-TDDFT data. However, the accuracy of this model depends on the ability of the parameterized single-valued coupling function to capture the complex rt-TDDFT data, and determining a suitable parameterization can be challenging.
Figure 1: (a) Illustration of a kinetic ion in a lattice, annotated with the radii used for the calculation of the ion’s neighbors. The outer radius $R_{out}$ portrays the cutoff of the UTTM parametrization $\alpha(\rho)$. The inner radius $R_{in}$ portrays the user-defined cutoff radius implemented in our approximation for the calculation of full friction components of the UTTM model in MDRANGE. (b) The energy loss, as a function of distance traveled, of a single 25 keV Ni ion in the $\langle 100\rangle$ center channel in Ni. (c) The predicted statistical range profile of 25 keV Ni ions along the $\langle 100\rangle$ channeling direction in Ni for varying $R_{in}$. The Ni $\alpha(\rho)$ parameterization cutoff is $5.0$ Å. Curves for $R_{in}=2.0-5.0$ Å coincide for the energy loss and range profiles.
III The Treatment of Trajectory-Dependent Electronic Energy Losses in MDRANGE
MDRANGE enables the efficient simulation of ion range profiles using MD through two key approximations. These approximations reduce computational complexity while maintaining accurate modeling of projectile ion dynamics, [8] but they also influence our implementation of the UTTM model:
1) The recoil interaction approximation (RIA): interatomic interactions between the material’s lattice nuclei are neglected. Instead, only the interactions between the projectile ion and the lattice nuclei are explicitly modeled, significantly simplifying the simulation complexity while capturing the essential projectile ion dynamics to high accuracy [8].
@@ III The Treatment of Trajectory-Dependent Electronic Energy Losses in MDRANGE
To enhance the computational efficiency of UTTM model calculations, we drew inspiration from the RIA. Specifically, we apply the calculated UTTM model tensorial friction force $\bm{\sigma}$ only to the projectile ion’s dynamics, neglecting friction effects on the lattice nuclei. This simplifies the treatment of the arising forces while preserving the accuracy in modeling ion trajectories.
The UTTM model was developed for use in full MD and accounts for correlated motion of lattice atoms; hence the calculations of the friction components include the positions and velocities of neighboring atoms in a large region around the projectile and their effects on each other. However, in MDRANGE, the target atoms are generated on-the-fly in perfect lattice positions with only thermal displacements perturbing the lattice. Since these atoms do not interact with each other due to the RIA approximation, and they move insignificantly during the time that the projectile passes the region, they can be treated in a mean-field approximation. To further optimize the UTTM model calculations in MDRANGE, and retain the small size of the lattice generated on-the-fly, we therefore defined an “inner radius” $R_{in}$ that includes a subset of the projectile ion’s closest neighbors. We illustrate this approximation in Fig. 1a. For neighbors within $R_{in}$, we account for their interactions with all their corresponding neighbors in the friction terms explicitly, including the projectile ion. For lattice nuclei outside $R_{in}$, but within the electron density cutoff $R_{out}$ for the UTTM model parametrization, we consider only the direct effect of the projectile ion, combined with a precalculated average contribution from other lattice atoms.
@@
Figure 1b shows that the energy losses calculated by the UTTM model for a single Ni projectile ion trajectory along the center $\langle 100\rangle$ channel are in complete agreement with the full UTTM model calculation ($R_{in}=R_{out}=5.0$ Å for the Ni parametrization) when $R_{in}$ is sufficiently large ($R_{in}>2.0$ Å in this case). As $R_{in}$ decreases, the energy losses of individual ions are progressively underestimated, due to the incomplete account of neighbor contributions. Consequently, the reduced ESP impacts the predicted statistical range profile, as shown in Fig. 1c.
@@
The correlated random forces defined in Equation 3 satisfy the fluctuation-dissipation theorem. In the MDRANGE implementation, the electronic temperature $T_{e}$ is not solved locally, and instead we set a constant temperature equal to the initial material temperature. We demonstrate in Fig. 2 that this does not significantly affect the range profile predictions in an exemplary case of 10 keV As ions implanted along the $\langle 100\rangle$ principal channeling direction in crystalline GaAs. The predicted range profiles do not show large discrepancies with different constant electronic system temperature values.
In Fig. 3, we further demonstrate that our approximations in the implementation of the UTTM model in MDRANGE do not affect the calculated energy losses, by comparing the energy losses predicted by MDRANGE with those obtained with LAMMPS, for a portion of the trajectory of a single 10 keV Si ion directed along the $\langle 100\rangle$ channel of Si, both in a center channel trajectory and for a trajectory shifted off-center along the direction of the channel. The UTTM implementation in LAMMPS offers full MD treatment of the projectile ion and lattice atoms and resolution of local electronic temperature, ensuring full energy conservation and satisfying the fluctuation-dissipation theorem.
@@
Figure 2: Integral range profiles of 10 keV As ions in GaAs along the $\langle 100\rangle$ principal channeling direction, using different constant values for the temperature of the electronic system. The range profiles were simulated the UTTM model for the ESP. We observe no significant differences in the resulting range profiles for different constant temperatures. The GaAs coupling function used here for the UTTM model was developed by Teunissen et al. [6].
Figure 3: Comparison of the rate of energy loss of 10 keV Si ions traveling along center-channel and off-channel $\langle 100\rangle$ trajectories in a pristine Si lattice, simulated using the UTTM model for the ESP with LAMMPS and with MDRANGE. In the off-channel trajectory, the projectile ion was shifted toward a neighboring lattice atom by $50\%$ of the distance between them in the center-channel trajectory. The predicted rates of energy loss show great agreement between both codes.

### v2

@@ I Introduction
Ion irradiation of materials is encountered in a wide range of research and industrial applications, including semiconductor doping [Williams1998], ion-beam lithography [lithography_review], and in the study of radiation-induced damage in nuclear materials [ranges_damage]. Understanding and predicting the effects of ion irradiation is crucial for optimizing material modification processes and for assessing irradiation-induced material damage [was2017]. Computational modeling provides a powerful and cost-effective approach for simulating ion-matter interactions and insight into these radiation effects.
@@
The predictive capability of computational models depends on the accurate description of interactions between energetic ions and target materials. These interactions lead to a reduction in the kinetic energy of the ion, quantified by the stopping power: $\frac{dE}{dx}=S_{e}+S_{n}$, where $S_{e}$ is the electronic stopping power (ESP) representing the energy transfer from the energetic ion to target electrons; and $S_{n}$ is the nuclear stopping power (NSP), which arises from interactions between energetic ions and target atomic nuclei [schmidt2013].
@@
Ion irradiation processes can be modeled using molecular dynamics (MD) simulations or Monte Carlo (MC) methods based on the binary collision approximation (BCA). MD simulations dynamically resolve atomic-scale interactions through Newtonian mechanics, enabling high-accuracy modeling of NSP using interatomic potentials [teunissen2023, sillanpaa2001]. In contrast, MC-based BCA methods model ion interactions and energy losses statistically, sacrificing atomic-scale accuracy in favor of computational efficiency, allowing for the simulation of many irradiation events. MDRANGE is an MD-based simulation code that was specifically developed to overcome the computational challenge of using MD to predict ion range profiles [nordlund_1994_mdrange], and offers an efficiency comparable to that of BCA methods with the atomic-scale accuracy of MD.
@@
The ESP is widely modeled as a scalar velocity-dependent retarding force according to Lindhard theory [lindhard1965], which has been expanded to improve accuracy across different energy regimes. Values for the ESP may be obtained numerically or from computational programs such as PSTAR for protons [PSTAR], MSTAR for heavier ions (3Li to 18Ar) [MSTAR], or SRIM (The Stopping and Range of Ions in Matter) for all ions. [SRIM] These programs combine theoretical frameworks, scaling laws, and empirical data to provide predictions for the average ESP of ions in materials.
@@
However, projectile ions can travel along trajectories in which they experience lower than average atomic and electronic densities. This effect is known as channeling and occurs when the ions are directed by interatomic potentials along the principal crystal axes (axial) or between planes or rows of atoms (planar) [lindhard1965, nordlund_2016_channel]. In these channeling geometries, projectile dynamics are affected less by lattice nuclei, and they lose their kinetic energy mainly through electronic interactions. Ions can also experience varying electronic interactions in these geometries because of the lower electron densities they encounter. These effects lead to a reduction in the total stopping power on average [sigmund1998] and result in higher ion ranges (penetration depths) in channeling, compared to non-channeling, geometries [nordlund_2016_channel].
@@
Real-time time-dependent density functional theory (rt-TDDFT) has been utilized to simulate the temporal evolution of quantum electronic interactions of energetic ions in materials [Nunez2025, lee_2020_tddft, lim_2016, tamm_2019_eph1, teunissen2023, jarrin_2021, ullah2018]. While rt-TDDFT can provide an accurate, first-principles description of electronic energy losses, its computational cost is prohibitive for large-scale ion irradiation simulations. In the approach outlined by Correa et al. [correa_2018_esfp], a trajectory- and energy-dependent scalar ESP can be obtained by analyzing the ion’s rate of energy loss as a function of its displacement in short trajectories and thus utilized in larger-scale simulations.
@@
Recently, Tamm et al. [tamm_2018_eph2] developed a model that uses rt-TDDFT energy loss data to parametrize a coupling function capturing the effects of electronic interactions on a projectile ion, dependent on its local environment, in a two-temperature molecular dynamics (TTMD) framework. This approach enables modeling of a trajectory-dependent ESP with ab initio accuracy in large-scale ion irradiation simulations. The model is implemented for the MD code LAMMPS [LAMMPS] in the USER-EPH plugin [tamm_2019_eph3].
@@
In this work, we present the implementation of the model developed by Tamm et al. [tamm_2018_eph2], which we hereby refer to as the unified two-temperature model (UTTM), in the MD-based code MDRANGE. This new implementation allows modeling ion ranges over micrometers in depth with ab initio based trajectory-dependent electronic energy losses, with statistics from tens of thousands of ions, which is inaccessible by full MD. Through comparison to other models in suitably chosen geometries, and to experimental measurements where such are available, this method also provides a tool for validating the parametrized coupling functions of the UTTM model. We investigate the performance of the UTTM model implementation in MDRANGE through simulations of ion range profiles in a variety of materials along both channeling and non-channeling trajectories.
@@ II Electronic Stopping Power Models
For any ESP model, the simulated ion range profiles vary along different crystal orientations, due to the different atomic environments encountered [nordlund_2016_channel]. Adjusting the ESP model further modifies these predictions. In this work, we compare trajectory-dependent ion range profile predictions in a variety of materials using different models for the ESP. We categorize two distinct ESP model types: (1) scalar models, whereby a velocity-dependent retarding force is applied to the projectile ion without accounting for the ion’s local environment, and (2) trajectory-dependent models, such as the UTTM model, which explicitly consider the ion’s local environment in the calculation of the retarding force on the projectile.
II.1 Scalar Electronic Stopping Power
The scalar model approach treats the ESP as a velocity-dependent friction force that decelerates the ion. In an MD framework, the force on an atom is defined by:
$\mathbf{f}_{I}=-\nabla_{I}U-\beta_{I}\left(v_{I}\right)\mathbf{v}_{I}$ (1)
@@ II.1 Scalar Electronic Stopping Power
where $U$ is the interatomic potential, and $\beta(v)$ is the scalar velocity-dependent ESP friction force. In this work, we employ values for the scalar ESP calculated by SRIM [SRIM], which is widely utilized in radiation damage and ion range studies. The ESP is calculated by SRIM semi-empirically, and is based on theoretical models fit to experimental data where it is available. This experimental data is scarce, particularly for heavier ions at low velocities. In these cases, the calculated ESP is extrapolated to the low-velocity regime, introducing uncertainties in modeling the electronic energy losses of slow ions [PREPRINT_Nickel].
@@
Moreover, a significant limitation of SRIM is that it does not account for the ordered lattice or electronic structure of materials. Instead, the ESP calculated by SRIM is generally valid for random ion trajectories in target materials. For trajectories where the local electronic and nuclear environments differ significantly from the average random trajectory, such as along channeling directions, electronic energy losses are not accurately modeled [sand2019].
@@
Unlike the semi-empirical approach, rt-TDDFT directly resolves the energy exchanged between projectile ions and the target electrons, which allows for a velocity- and trajectory-dependent description of electronic energy losses. A scalar ESP force on the ion can be calculated for chosen trajectories as outlined by Correa et al. [correa_2018_esfp]. While this provides higher fidelity in modeling the ESP of ions which travel along channels, it does not account for the local ion environment. Consequently, it does not accurately model electronic energy losses for projectiles which are either not significantly channeled, or have been scattered out of the channel. Nevertheless, the electronic energy losses for strongly channeled ions are captured correctly in range simulations employing a constant rt-TDDFT value for the ESP. Hence, this approach yields accurate predictions of the maximum ion range $R_{max}$ [sand2019], which is typically determined in ion range experiments [eriksson1967].
II.2 Trajectory-Dependent Electronic Stopping Power
@@
Accurately capturing the ESP requires a model which accounts for variations in the local electronic environment encountered along ion trajectories. The trajectory-dependent UTTM model is based on Langevin dynamics, in which particles in a system experience frictional forces and corresponding random forces via the fluctuation-dissipation theorem. The UTTM model modifies the traditional Langevin dynamics by replacing forces on individual atoms with correlated many-body forces, and a tensorial frictional force describes the ESP acting on the atoms [tamm_2018_eph2]. The force on an atom is defined by:
@@
where $\rho_{I}(r_{IJ})$ is the contribution to the electron density of atom $I$ at a distance $r_{IJ}$, and $\bar{\rho}_{J}=\sum_{I\neq J}\rho_{I}(r_{IJ})$ is the total local electron density experienced by atom $J$. The vector $\bm{e}_{IJ}$ is the unit vector joining atoms $J$ and $I$. The coupling function $\alpha_{J}=\alpha(\bar{\rho}_{J})$ is dependent on the total electron density $\bar{\rho}_{J}$ experienced by atom $J$. The correlated random forces are determined by a set of independent white noise Gaussian variables $\langle\xi_{I}(t)\rangle_{I}$ given by:
@@
where $T_{e}$ is the temperature of the electronic system [tamm_2018_eph2]. The $\alpha(\bar{\rho})$ coupling component is determined from ab initio rt-TDDFT data. However, the accuracy of this model depends on the ability of the parameterized single-valued coupling function to capture the complex rt-TDDFT data, and determining a suitable parameterization can be challenging.
III Simulation Methodology
III.1 MDRANGE Approximations
MDRANGE enables the efficient simulation of ion range profiles through two key approximations. These approximations reduce computational complexity while maintaining accurate modeling of projectile ion dynamics[nordlund_1994_mdrange]:
1) The recoil interaction approximation (RIA): interatomic interactions between the material’s lattice nuclei are neglected. Instead, only the interactions between the projectile ion and the lattice nuclei are explicitly modeled, significantly simplifying the simulation complexity while capturing the essential projectile ion dynamics to high accuracy [nordlund_1994_mdrange].
@@ III.1 MDRANGE Approximations
III.2 Ion Range Profile Calculation
Projectile ions are initialized at randomly distributed positions near the surface of the target. The target materials are thermalized with displacements determined by the Debye model, as detailed in Ref. [debye_disp]. Nuclear interactions between projectile ions and lattice atoms are modeled using the Ziegler–Biersack–Littmark (ZBL) repulsive potential [ZBL_book]. Each ion trajectory is simulated until the kinetic energy of the projectile reaches 1 eV. In accordance with ion range experiments [eriksson1967], we present the ion range profiles as integrated distributions which show the fraction of implanted ions that reached a given range.
III.3 Implementation of Trajectory-Dependent Electronic Energy Losses in MDRANGE
We have implemented the UTTM model in MDRANGE, to incorporate trajectory-dependent electronic energy losses in ion range simulations. The UTTM model was originally developed for use in full MD and accounts for correlated motion of lattice atoms, hence the calculations of the friction components include the positions and velocities of neighboring atoms in a large region around the projectile and their effects on each other.
Analogous to the RIA, in our implementation of the UTTM model in MDRANGE we apply the UTTM tensorial forces only to the projectile ion’s dynamics, neglecting electronic effects on the lattice nuclei. This simplifies the treatment of the arising forces while preserving the accuracy in modeling ion trajectories. In Fig. 1, we demonstrate that this approximation does not affect the calculated energy losses through comparison with the USER-EPH package in LAMMPS [tamm_2018_eph2, tamm_2019_eph3], which offers full MD treatment of the projectile ion and lattice atoms. We compare portions of single 10 keV Si ion trajectories directed along the $\langle 100\rangle$ channel in Si, both in a center channel trajectory and for a trajectory shifted off-center along the direction of the channel. For these short trajectories, the maximum pointwise difference of MDRANGE relative to LAMMPS is less than 0.03%.
Figure 1: Comparison of the rate of energy loss of 10 keV Si ions traveling along center-channel and off-channel $\langle 100\rangle$ trajectories in an Si lattice, simulated using LAMMPS and MDRANGE. The ESP was modeled using the UTTM model with a quadratic $\alpha(\bar{\rho})$ function developed by Jarrin et al. [jarrin_2021]. The target was initialized at 0 K without thermal displacements. In the off-channel trajectory, the projectile ion was shifted toward a neighboring lattice atom by $50\%$ of the distance between them in the center-channel trajectory.
Because target lattice atoms do not interact with each other due to the RIA, they move insignificantly during the time that the projectile passes the region, and they can be treated in a mean-field approximation. To further optimize the UTTM model calculations in MDRANGE, and retain the small size of the lattice generated on-the-fly, we therefore defined an “inner radius” $R_{in}$ that includes a subset of the projectile ion’s closest neighbors. We illustrate this approximation in Fig. 2a. For neighbors within $R_{in}$, we account for their interactions with all their corresponding neighbors in the friction terms explicitly, including the projectile ion. For lattice nuclei outside $R_{in}$, but within the electron density cutoff $R_{out}$ for the UTTM model parametrization, we consider only the direct effect of the projectile ion, combined with a precalculated average contribution from other lattice atoms.
@@ III.3 Implementation of Trajectory-Dependent Electronic Energy Losses in MDRANGE
Figure 2b shows that the energy losses calculated by the UTTM model for a single Ni projectile ion trajectory along the center $\langle 100\rangle$ channel are in complete agreement with the full UTTM model calculation ($R_{in}=R_{out}=5.0$ Å for the Ni parametrization) when $R_{in}$ is sufficiently large ($R_{in}\geq 3.0$ Å in this case). As $R_{in}$ decreases, the energy losses of individual ions are progressively underestimated, due to the incomplete account of neighbor contributions. Consequently, the reduced ESP impacts the predicted integrated range profile, as shown in Fig. 2c.
Figure 2: (a) Illustration of a kinetic ion in a lattice, annotated with the radii used for the calculation of the ion’s neighbors. The outer radius $R_{out}$ portrays the cutoff of the UTTM parametrization $\alpha(\bar{\rho})$. The inner radius $R_{in}$ portrays the user-defined cutoff radius implemented in our approximation for the calculation of full friction components of the UTTM model in MDRANGE. (b) The energy loss, as a function of distance traveled, of a single 25 keV Ni ion in the $\langle 100\rangle$ center channel in Ni without thermal displacements. (c) Calculated integral range profile of 25 keV Ni ions along the $\langle 100\rangle$ channeling direction in Ni, thermalized at 300 K, for varying $R_{in}$. Ranges were calculated using the method described in Sec. III.2 for 50,000 ion implants. We employed the Ni $\alpha(\bar{\rho})$ parameterization developed by Caro et al. [tamm_2019_eph1], which has a cutoff of $5.0$ Å. Curves for $R_{in}=3.0$ and $5.0$ Å coincide for the energy loss and range profile.
@@
The electronic temperature $T_{e}$ which appears in Equation 4 is not solved locally in MDRANGE, and instead we set a constant temperature equal to the initial material temperature. We demonstrate in Fig. 3 that this does not significantly affect the range profile predictions in an exemplary case of 10 keV As ions implanted along the $\langle 100\rangle$ principal channeling direction in crystalline GaAs. The predicted range profiles do not show large discrepancies with different constant electronic system temperature values.
@@
Figure 3: Integral range profiles of 10 keV As ions in GaAs along the $\langle 100\rangle$ principal channeling direction using different constant values for the temperature of the electronic system. The atomic system was thermalized to 300 K in all cases. The range profiles were calculated using the method described in Sec. III.2 for 100,000 ion implants, with the UTTM model for the ESP. The GaAs coupling function used here for the UTTM model was developed by Teunissen et al. [teunissen2023].
Figure 4: Calculated integral ion range profiles of 10 keV Si ions in diamond Si along the (a) $\langle 100\rangle$, (b) $\langle 111\rangle$, and (c) $\langle 110\rangle$ channeling directions, and (d) along a random direction, using different models for the ESP. Each channeling plot contains the range profiles predicted using the UTTM model with two parametrizations, both produced by Jarrin et al. [jarrin_2021], which differ in the functional form of the $\alpha(\bar{\rho})$ coupling function. Range profile predictions using the scalar ESP calculated using SRIM are plotted for all directions, and using scalar TDDFT calculated by Nuñez-Palacio et al. [Nunez2025] for the channeling directions. In the $\langle 100\rangle$ and $\langle 110\rangle$ directions, the stopping corresponds to the center-channel, and in the $\langle 111\rangle$ direction it corresponds to the “half-center” channel. Note the different axes scales, particularly the logarithmic x-axis scale in (c).